40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Message ID: 946     Entry time: Wed Aug 18 12:03:23 2010
Author: Alastair 
Type: Electronics 
Category: GYRO 
Subject: PDH boxes 

I was looking at the error signal from the two loops, wondering why the CW one looked so small.  I decided to swap around the locking loops  as we discussed at the meeting the other day.

To begin with I swapped the laser feedback onto the feedback signal from the CW loop.  It would barely lock, so I tried sweeping the laser freq to see what the error signal looked like.  It looked very small.

I decided to swap the PDH boxes to see what that did.  The signal was now very large.  I went back to locking the laser to the cavity using the CCW loop, and took two error signal measurements, one using each PDH box, with exactly the same settings and connections.  You can see from the pictures below that they are very different.  The one with the really small signal is the one we were using to lock the AOM before.

I will take the PDH boxes down to investigate.

**EDIT** Internally the two boxes don't appear to be the same though I'm not sure how big the differences are.  Firstly, the working one has a 1.9MHz low pass filter t-d to a 50ohm terminator going from the output of the mixer to the 'mixer out' bnc connector.  The second box has only a 5MHz low pass filter there.  I swapped the parts out of the first box into the second to test if there was any difference and there was not.

Second obvious difference is that the working box has a 10K resistor added going from the center pin of the piezo sweep input to the op-amp side of R19 the piezo driver output amp.  It would seem that this is basically just bypassing the relay that the piezo sweep switch uses. 

**EDIT 2** I found Rana's elog entry that tracks the changes to box 1 (the working box).  It is here: http://nodus.ligo.caltech.edu:8080/AdhikariLab/642

I don't see any mention that any work has taken place on the second box.  I'll print the schematic and mark up the changes that were made to box 1 on it.

Attachment 1: TEK00000.PNG  15 kB  | Hide | Hide all
TEK00000.PNG
Attachment 2: TEK00005.PNG  14 kB  | Hide | Hide all
TEK00005.PNG
ELOG V3.1.3-